OpenStudy (anonymous):

how to simplify this (1^-4 + 2^-4 + 3^-4 + 4^-4 + ...)/(1^-4 + 3^-4 + 5^-4 + 7^-4 + ...)

OpenStudy (raden):

let p = 1^-4 + 2^-4 + 3^-4 + 4^-4 + ... q = 1^-4 + 3^-4 + 5^-4 + 7^-4 + ... see your question is p/q. now, we get the value of p - q = 2^-4 + 4^-4 + 6^-4 + 8^-4 + 10^-4 + ... or p - q = 2^-4(1^-4 + 2^-4 + 3^-4 + 4^-4 + ... ) p - q = 2^-4(p) p - q = p/16 p - p/16 = q (15/16) p = q p/q = 16/15 so, it is 16/15

OpenStudy (anonymous):

still confused :(

OpenStudy (anonymous):

@RadEn 's argument is this You can define your two variables, p and q, to be the numerator and the denominator of the expression. So you have \[ p = 1^{-4} +2^{-4} + 3^{-4} +4^{-4} ..... = \sum_{n=1}^{\infty} n^{-4}\] \[ q = 1^{-4} + 3^{-4} + 5^{-4} + 7^{-4}..... = \sum_{n=1}^\infty (2n-1)^{-4}\] if you subtract those from each other, all of the odd bases cancel out \[ \color{purple}p-q = ( \cancel{\color{purple}{1^{-4}} - 1^{-4}}) +\color{purple} {\textbf 2^{-4}} + (\cancel{\color{purple}{3^{-4}} - 3^{-4}} )+ \color{purple}{\textbf 4^{-4}} + (\cancel{\color{purple}{5^{-4}} - 5^{-4}}) + \color{purple}{\textbf 6^{-4}}..... \] leaving you with \[ p -q = 2^{-4}+4^{-4}+6^{-4}+8^{-4}...\] from which you can factor out the 2^-4 \[ p - q = 2^{-4}(1^{-4}+2^{-4}+3^{-4}...) = 2^{-4}(p)\] solve for q in terms of p \[ q = p (1-\frac{1}{16})\] Seeing then that \[ \frac{p}{q} = \frac{16}{15} \]

OpenStudy (raden):

thanks, @AllTehMaffs . i think that's make more sense now :)

OpenStudy (anonymous):

It's a nice proof you found :)

OpenStudy (raden):

but i cant explain it more, actually my englsih so bad. :)

OpenStudy (raden):

* english

OpenStudy (anonymous):

Your maths are awesome! So no worries ^^

OpenStudy (raden):

moreover i cant using LATEX, if ican do like you can make be happy to see it your writting is beautiful

OpenStudy (anonymous):

:) thank you :)

OpenStudy (anonymous):

thanks guys, that's make more understand now